# Block-corrected Modularity for Community Detection

### v2

@@ II Block-corrected modularity
In the following, we focus on unweighted, directed networks characterized by their asymmetric adjacency matrix $\mathbf{A}$. Element $A_{ij}$ of the adjacency matrix $\mathbf{A}$ equals one if there is an edge from node $i$ to node $j$ and zero otherwise. $V$ is the set of nodes in the network. The number of edges and nodes in the network is denoted by $m$ and $|V|$, respectively.
@@ III.2.2 Temporal connectivity pattern in null models
Figure 3(b) depicts the connectivity matrix $\mathbf{M}$ of a network drawn from the skewed network model with two unknown communities, along with connectivity matrices of the three null models. The effect of the known attribute (time) is encoded in the $2\times 2$ blocks of nonzero probability of connectivity near the diagonal representing adjacent time layers, as well as the $2\times 2$ block at the corner representing connections between the first and the last layer. The structure induced by the unknown attribute is encoded in the difference of colors between diagonal and anti-diagonal elements in each $2\times 2$ block.
@@ III.2.3 Susceptibility to erroneous partitioning
Figure 4: Difference between modularity values for the erroneous candidate partitions, where the first and last $1/4$ of temporal blocks are placed in one community and the rest in another, and that of the true partition over an ensemble of random networks with varying average in-group degrees. All networks are realizations of the skewed model (Eq. (10)), with $t_{\mathrm{max}}=12$ temporal blocks and $N=100$ nodes per block. $Q_{\mathrm{true}}>Q_{\mathrm{cp}}$ indicates that a modularity prefers the true partition to the erroneous candidate. Error bars show 95% confidence intervals.
@@ IV.1 Spectral partitioning
Given a network, finding two communities is equivalent to discovering a membership vector, $s$, whose elements are community labels of $+1$ and $-1$ of all nodes, $\mathbf{s}=(s_{1},\ldots,s_{|V|})^{\intercal}$, $s_{i}\in\{+1,-1\}$. Modularity value in Eq. (1) can be thus re-written as a function of membership vector $\mathbf{s}$ and modularity matrix $\mathbf{B}$, whose element $B_{ij}$ is equal to $A_{ij}-P_{ij}$, and
@@
where $\tilde{\mathbf{s}}$ is the membership vector of community $C$ to be solved, and $\tilde{\mathbf{B}}^{(C)}$ a generalized modularity matrix of size $|C|\times|C|$ where $\tilde{B}_{ij}^{(C)}=B_{ij}-\delta_{ij}\sum_{k\in C}B_{ik}$ for $i,j\in C$ [14, 7]. From this point forward the superscript $(C)$ for a matrix denotes a submatrix containing rows and columns in $C$. The bipartition of community $C$ is executed only when the corresponding $\Delta Q$ is positive, indicating that performing the bipartitioning results in a gain in modularity value. The iterative bipartitioning halts when modularity value $Q$ stops increasing, meaning $\Delta Q$ is non-positive. In the very first step of the algorithm, community $C$ corresponds to the entire network ($C=V$) in which case, assuming no self-loops are present, $\tilde{\mathbf{B}}^{(V)}=\mathbf{B}$ and $\tilde{\mathbf{s}}=\mathbf{s}$, resulting in $\Delta Q$ in Eq. (16) being equal to $Q$ in Eq. (15).
@@
(((( $\mathbf{P}\mathbf{x})_{i}=\sum_{j=1}^{n}P_{ij}x_{j}=\sum_{j=1}^{n}\kappa_{j}^{\mathrm{in}}\kappa_{i}^{\mathrm{out}}L_{t_{i}t_{j}}x_{j}$
@@ Appendix A Computing modularity values across known and unknown attributes
Under the block null model as defined in Eq. (2), the term $k_{i}^{\mathrm{out}}/K_{x_{i}}^{\mathrm{out}}$ in $\tilde{p}_{ij}$ becomes $2/|V|$ since there are $|V|/2$ nodes in community $x_{i}$, the known attribute. Thus $\tilde{p}_{ij}$ transforms into $4/|V|^{2}L_{x_{i}x_{j}}$ where $L_{x_{i}x_{j}}$ is the number of edges between nodes with label $x_{i}$ and nodes with label $x_{j}$.
We assume edge densities in each community of $x=0$ and $x=1$ are equal, as well as between the two, that is, $L_{00}=L_{11}$ and $L_{01}=L_{10}$. Node $i$ can be connected to a node whose label for $y$ could be the same with probability $p_{1}^{y}$, or different with probability $p_{0}^{y}$. There are $|V|/4-1$ possible nodes tagged with the same $y$ as node $i$ and $|V|/4$ with a different value for $y$. Since there are $|V|/2$ possible nodes with the same value for $x$ as node $i$ we get
@@
Elements $L_{01}=L_{10}$ of matrix $\mathbf{L}$ are equal to $|V|/2(|V|p_{0}^{x}/4+p_{1}^{y}+|V|p_{0}^{x}p_{0}^{y}/4)=|V|^{2}p_{0}^{x}(p_{0}^{y}+p_{1}^{y})/8$ since there are $|V|/2$ nodes in each community with respect to $x$ and for every node the destination node is in the same community with respect to $y$ with probability $p_{1}^{y}$ and probability $p_{0}^{y}$ of being in different communities with respect to $y$.
@@
We substitute the right-hand side of Eq. (27) in Eq. (5) and plug in $42(L_{00}+L_{01})=|V|^{2}(p_{1}^{x}+p_{0}^{x})(p_{1}^{y}+p_{0}^{y})/$ 4 for $m$ to arrive at
@@
are held, the modularity maximization method using the proposed block-corrected modularity will successfully recover hidden attribute $y$, but does not continue to needlessly split the found communities based on $(x,y)$. Since $Q^{\mathrm{bc}}_{x}\approx 0$, the first part of the inequality simplifies into $p^{y}_{1}>p^{y}_{0}$, i.e., that the hidden attribute must be assortative. The second inequality translates to
@@
which leads to $p_{1}^{x}(3p_{0}^{y}-p_{1}^{y})>p_{0}^{x}(p_{1}^{y}+p_{0}^{y})$. Since the left-hand side is positive, this implies that $03p_{0}^{y}-p_{1}^{y}>$ 0 should hold and
@@
For example, with $p_{1}^{x}=0.9$, $p_{0}^{x}=0.05$, $p_{1}^{y}=0.35$ and $p_{0}^{y}=0.25$, we get $Q^{\mathrm{dir}}_{x}=0.45$, $Q^{\mathrm{bc}}_{x}\approx 0$, $Q^{\mathrm{dir}}_{y}=Q^{\mathrm{bc}}_{y}=0.084$, $Q^{\mathrm{dir}}_{xy}=0.30$ and $Q^{\mathrm{bc}}_{xy}=0.078$ which satisfies both inequalities above. In this scenario, directed modularity splits the network into two communities based on the known attribute $x$ but does not proceed to split it again into four communities based on $(x,y)$, while block-corrected modularity correctly identifies the unknown attribute $y$ and does not further divide the network.

### v3

@@ II Block-corrected modularity
[…]
@@ III.2.2 Temporal connectivity pattern in null models
Figure 3(b) depicts the connectivity matrix $\mathbf{M}$ of a network drawn from the skewed network model with two unknown communities, along with connectivity matrices of the three null models. The effect of the known attribute (time) is encoded in the $\times 2$ 2 blocks of nonzero probability of connectivity near the diagonal representing adjacent time layers, as well as the $\times 2$ 2 block at the corner representing connections between the first and the last layer. The structure induced by the unknown attribute is encoded in the difference of colors between diagonal and anti-diagonal elements in each $\times 2$ 2 block.
@@ III.2.3 Susceptibility to erroneous partitioning
Figure 4: Difference between modularity values for the erroneous candidate partitions, where the first and last 1/41/41 / 4 of temporal blocks are placed in one community and the rest in another, and that of the true partition over an ensemble of random networks with varying average in-group degrees. All networks are realizations of the skewed model (Eq. (10)), with $t_{\mathrm{max}}=12$ temporal blocks and $N=100$ nodes per block. $Q_{\mathrm{true}}>Q_{\mathrm{cp}}$ indicates that a modularity prefers the true partition to the erroneous candidate. Error bars show 95% confidence intervals.
@@ IV.1 Spectral partitioning
Given a network, finding two communities is equivalent to discovering a membership vector, $s$, whose elements are community labels of +1+1+ 1 and 1-1- 1 of all nodes, $\mathbf{s}=(s_{1},\ldots,s_{|V|})^{\intercal}$, $s_{i}\in\{+1,-1\}$. Modularity value in Eq. (1) can be thus re-written as a function of membership vector $\mathbf{s}$ and modularity matrix $\mathbf{B}$, whose element $B_{ij}$ is equal to $A_{ij}-P_{ij}$, and
@@
[…]
@@
((( $\mathbf{P}\mathbf{x})_{i}=\sum_{j=1}^{n}P_{ij}x_{j}=\sum_{j=1}^{n}\kappa_{j}^{\mathrm{in}}\kappa_{i}^{\mathrm{out}}L_{t_{i}t_{j}}x_{j}$
@@ Appendix A Computing modularity values across known and unknown attributes
[…]
@@
[…]
@@
We substitute the right-hand side of Eq. (27) in Eq. (5) and plug in $2(L_{00}+L_{01})=|V|^{2}(p_{1}^{x}+p_{0}^{x})(p_{1}^{y}+p_{0}^{y})/4$ for $m$ to arrive at
@@
[…]
@@
which leads to $p_{1}^{x}(3p_{0}^{y}-p_{1}^{y})>p_{0}^{x}(p_{1}^{y}+p_{0}^{y})$. Since the left-hand side is positive, this implies that $3p_{0}^{y}-p_{1}^{y}>0$ should hold and
@@
[…]